# CoSQA+: Enhancing Code Search Evaluation with a Multi-Choice Benchmark and Test-Driven Agents

### v5

@@ Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 92.0%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
Index Terms:
software engineering, information search and retrieval, human-computer interaction.
@@ I Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development [1, 2, 3, 4, 5]. Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6]. Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12]. According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
However, existing approaches face significant challenges in scalability, annotation quality, and evaluation metrics:
@@
    Accuracy Limitations in Annotation Methods. Both human and LLM-based annotation approaches suffer from accuracy issues. Human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. While Li et al. leverage LLMs to address scalability issues [9], these models struggle to reliably determine whether code functionally satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. This fundamental limitation affects the quality of both human and LLM-based annotations [13].
@@
    Inadequate Evaluation Metrics for Code Search. Most code search benchmarks employ metrics like MRR or Pass@k [10] which primarily evaluate single-result retrieval performance. This overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k [7] exist for evaluating multiple code matches, they underemphasize the importance of top-ranked results. This is particularly problematic as our survey reveals users heavily rely on first-page search results, highlighting the need for better evaluation metrics.
To address these limitations, we present a new benchmark CoSQA+, with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks [14]. We use it to measure 8 code search methods, 7 of which are embedding-based.
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 400 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 92.0%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
@@
  • To the best of our knowledge, we develop the first Multiagent annotator with 92.0% accuracy for code search dataset.
@@
  • We publicly release CoSQA+ to facilitate further research in semantic code search.
@@ II Motivation Survey
Do developers really need multi-choice code search? To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.Fig. 1 provides a multi-choice code search example
@@
The survey targets 102 participants, predominantly undergraduate students (89.22%) with a smaller proportion of master’s students (8.82%). Most of the participants (92.16%) have formal computer science education, and their experience in Python programming varies. The majority (66.67%) have between 1 and 5 years of programming experience, while 30.39% have less than one year of experience. The participants are selected from programmers using Python as their primary programming language (52.94%). The survey is conducted anonymously through an online platform. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
@@
Our survey reveals several significant patterns in code search behavior. Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution. This suggests that the current code search task, which typically prioritizes a single result, may need to evolve to provide multiple relevant code snippets. Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%). These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better. When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
@@ IV-A Query and Code Collection
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes [8]. We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57]. (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process [58]. By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
@@
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7]. According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability [59]. To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
@@ IV-B Candidate Pairs Construction
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. Fig.  3 shows a detailed breakdown of query distributions by the number of exactly matched codes.
@@
The relatedness between the query and code is quantified by the cosine similarity, calculated as:
@@ IV-C Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [60], using the algorithm shown in Algorithm 1. Initially, A preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation.
@@
[…]
@@
10:     while $exec\_res.\text{error}\ \textbf{and}{}\ try\_time\leq max\_try\_time$ do
@@
[…]
@@
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, we generate the test program based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
@@ V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
@@
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG [14] and MRR [12]. NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
$NDCG@k=\frac{DCG@k}{IDCG@k}$
where DCG@k is the Discounted Cumulative Gain up to position k, and IDCG@k is the ideal DCG@k (the maximum possible DCG@k). The final NDCG score is averaged across all queries. MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
@@
For the second experiment, We fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with NDCG@10 and MRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
@@
After the volunteers compare the two datasets, we come to the result that out of the 2000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.

### v6

@@ Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 93.9%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We publicly release both CoSQA+_all, which contains 412,080 agent-annotated pairs, and CoSQA+_verified, which contains 1,000 human-verified pairs, at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
@@ I Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development [1, 2, 3, 4, 5]. Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6]. For the given query, while existing semantic code search studies focus on retrieving relevant codes, we define a different task: to retrieve functionally matching codes that satisfy the query’s requirements when executed in a complete context.
Although one query can be matched by multiple codes, existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12]. This design leads to a critical mismatch with real-world practices. According to our survey of 200 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 7.79 times per programming day on average. For each query, they typically reference 2.83 code examples to find the necessary information. Furthermore, participants report that 63.2% of their queries commonly result in multiple valid code snippets. This reveals that developers operate in a multi-choice context, yet existing benchmarks adhere to a single-choice evaluation paradigm with only one "correct" code snippet per query.
The construction of benchmarks that reflect this multi-choice reality faces significant challenges, as methodologies designed for the prevalent one-to-one task (e.g., CoSQA [8], CodeSearchNet [7]) are hampered by issues of scalability, accuracy, and evaluation:
@@
    Accuracy Limitations in Annotation Methods. Both human and LLM-based annotation approaches suffer from accuracy issues. Human annotators typically assess code matching based on surface-level comprehension rather than verifying functional correctness [13] through test execution, which may lead to inaccurate annotations due to a lack of full understanding of syntax, semantics, and behavioral correctness. While Li et al. leverage LLMs to address scalability issues [9], these models struggle to reliably determine whether code functionally satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. This fundamental limitation affects the quality of both human and LLM-based annotations [14].
@@
    Lack of Evaluation in Multi-Choice Scenarios. One-to-one Conventional one-to-one matching benchmarks typically rely on MRR, which measures the rank position of the first relevant code snippet. However, in practical scenarios, developers often seek multiple code examples for reference, where the key concern is how many relevant snippets appear within the top-10 results or on the first page. MRR and other single-choice benchmarking approaches fail to capture this dimension of performance.
To address these limitations, we present a new benchmark CoSQA+ , comprising two releases: CoSQA+_all with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, and CoSQA+_verified with 1,000 human-verified pairs that serve as a gold-standard evaluation subset. The benchmark is specifically designed for multi-choice code search and adopts MAP as the primary evaluation metric.
To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Specifically, exactly matched codes refer to codes whose functionality fully satisfies or exceeds the requirements specified in the query. Furthermore, we evaluate 8 code search methods—7 of which are embedding-based.
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 1000 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 93.9%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both MAP(Mean Average Precision)@10 and MRR metrics. Furthermore, we demonstrate that the test-driven agent generalizes beyond Python, achieving consistent accuracy gains of 10.33% on average across PHP, Java, and Go (RQ5).
@@
  • To the best of our knowledge, we develop the first Multiagent annotator with 93.9% accuracy for code search dataset.
@@
  • We publicly release both CoSQA+_all and CoSQA+_verified to facilitate further research in semantic code search.
@@ II Motivation Survey
Do developers really need multi-choice code search? To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process. Fig. 1 provides a multi-choice code search example. This example illustrates a common real-world scenario where query ambiguity leads to multiple legitimate interpretations. The term "illegal character" can reasonably refer to filesystem-invalid characters, non-ASCII characters, escape sequences, or special symbols, each representing valid use cases in different programming contexts.
@@
We recruited participants through online anonymous questionnaire platforms, targeting two main groups: current university students and working technical personnel in computer science or related fields. The survey collected responses from 200 participants, including 148 current students and 52 working professionals. In terms of educational background, the majority of participants hold bachelor’s degrees (87.5%) with a smaller proportion holding master’s or doctoral degrees (11.5%). Their Python programming experience varies, with the majority (59.0%) having between 1 and 5 years of programming experience, while 16.0% have between 5 and 10 years of experience. Among all participants, 64.0% use Python as their primary programming language. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
@@
Our survey reveals several significant patterns in code search behavior across different user groups. Importantly, the need for multiple code examples is not limited to undergraduate students but extends to experienced developers as well. We find that 75.0% of working technical professionals typically require 2-3 code examples to understand and implement a solution, compared to 63.5% of current students. This finding challenges the assumption that multi-choice code search primarily serves educational purposes for beginners. In fact, experienced developers demonstrate an even higher tendency to reference multiple implementations, particularly when dealing with complex or unfamiliar tasks. Additionally, the participants reported that 63.2% of their queries commonly result in multiple code snippets that meet their requirements, indicating that query ambiguity leading to multiple valid solutions is a prevalent phenomenon in real-world development scenarios. Code searchers prefer function codes with clear input and output (85.5%), indicating a strong preference for code examples that explicitly define parameters and expected results regardless of their experience level. Participants evaluate search results based on several key factors, primarily focusing on solution differences (66.5%) and implementation relevance (66.5%). These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better. When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
@@ IV-A Query and Code Collection
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes [8]. We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57]. (2) Through heuristic filtering and manual annotation, the queries maintain authentic code search intent while reducing non-code-search-related noise. This careful curation helps maintain high data quality and reduces noise in the training process [58]. By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
@@
To the best of our knowledge, CodeSearchNet is the only large-scale, real-world Python function dataset explicitly designed for code search. While other benchmarks (e.g., CoSQA [8], FunSearch [59], AdvTest [10]) include Python code, their codebases are derived from CSN — either by reusing its functions directly or by further processing its splits. This makes CSN the de facto origin for Python code-search evaluation. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7].
According to our survey, 85.5% of developers report that functions with explicitly defined inputs and outputs best meet their needs, reflecting real-world developer preferences. Such functions are consistently rated higher in readability, testability, and maintainability [60]. While functions with outputs but no inputs may still be relevant and testable in some contexts, they are often considered poorly designed. These functions typically exhibit unclear logic, rely on implicit global states or hard-coded constants, and are difficult to reuse across different scenarios [60, 61].
To assess data quality and relevance, we carefully reviewed 1,995 functions and found that only 16.25% of functions without input or output parameters are relevant to typical code search tasks, compared to 80.97% of functions with both input and output parameters. Including parameter-less functions would therefore introduce noise and degrade overall dataset quality.
Therefore, to construct a high-quality Python code corpus from CodeSearchNet’s 1,156,085 functions, we employ a rule-based AST filter to retain only those functions having at least one input parameter and a valid return statement, yielding 653,994 candidates. By concentrating on these clearly specified and testable functions, our dataset preserves sufficient scale, higher quality, and provides a solid foundation for test-driven agents to design and execute meaningful test programs, ultimately improving downstream model quality and supporting accurate code search evaluation.
@@ IV-B Candidate Pairs Construction
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matched codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Cosine similarity is calculated as:
Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. Fig.  3 shows a detailed breakdown of query distributions by the number of exactly matched codes.
@@
The relatedness between the query and code is quantified by the cosine similarity.
@@ IV-C Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [62], using the algorithm shown in Algorithm 1. Initially, a preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear. For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation.
@@
4:else if $assessment=\text{``clearly non-match''}$ then
@@
[…]
@@
16:  return $\text{final\_arbiter}(q,c,test\_prog,exec\_res)$
@@
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, we generate the test program based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation. To be clear, the mere fact that code runs does not guarantee it is relevant to the query. The agents, therefore, generate test programs that execute the code and compare its actual outputs or side effects with the results expected from the query’s intent. Execution logs and assertion outcomes are then passed to the final arbiter, which makes a holistic judgment rather than relying solely on execution success or failure. In this way, verification shifts from a literal understanding of the snippet to functional validation of its behaviour.
@@ V Evaluation
RQ5: Can the test-driven agent generalize across different programming languages?
@@ V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA by examining how well models can learn from these datasets. Here, we treat CoSQA and CoSQA+ as training datasets rather than benchmark suites, focusing on their effectiveness for model training rather than their utility as evaluation benchmarks. Note that this dual functionality — serving as both training corpus and evaluation benchmark — follows established precedent in code search research, where datasets like CodeSearchNet [7], CoSQA [8] and StaQC [11] similarly function in both capacities. For this specific comparison, we focus on the training aspect to assess dataset quality and learning effectiveness.
@@
Following previous work, we employ two metrics to evaluate the performance of code search models: MAP and MRR [12]. MAP measures the overall ranking quality by averaging precision scores across all relevant results for each query. For a given query $q_{i}$ with $R_{i}$ relevant results, we calculate Average Precision (AP) as:
$AP@k=\frac{1}{R_{i}}\sum_{j=1}^{k}P(j)\cdot rel(j)$
where $P(j)$ is the precision at position $j$, and $rel(j)$ is an indicator function that equals 1 if the result at position $j$ is relevant. The final MAP score is obtained by averaging AP across all queries:
$MAP=\frac{1}{N}\sum_{i=1}^{N}AP@k_{i}$
where $N$ is the total number of queries. This metric is particularly suitable for code search as it properly handles cases where multiple correct results may exist for a query.
MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
@@
For the second experiment, We fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with MAP and MRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
@@
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in MAP@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.